# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v2

@@ IV Neuromorphic Applications
IV-A Stable and Unstable Memristors
@@
We may consider two types of memristor: those which are highly unstable and those that are stable. For the model as presented in Section III, the stability of the memristor can be characterised by the rate of attainment of its equilibrium point, which is determined in large part by the magnitude of the energy barrier $E_{a}$. Unstable devices may be useful for the emulation of neuronal body dynamics, where the interplay between volatility and the application of a stimulating input (for example, current spikes) can result in self-limiting spiking responses. In this context, resistive drift happening over short timescales is analogous to the leakage of membrane potential. In contrast, stable memristors may be more useful for the emulation of synaptic weights. In this context, we can model a combination of short and long term memory effects: Short Term Potentiation (STP) can arise for the programming of a stable memristor to a particular state while the volatility is high. The volatility may cause the state variable to partially revert to its equilibrium point, but eventually dissipates, resulting in stability of the resistive state ( Long Term Potentiation (LTP)). For stable memristors, resistive drift is analogous to the phenomenon of synaptic weight decay.
IV-B Volatility and its Impact
@@
The inclusion of volatility in a memristive model adds a key characteristic that could prove crucial to the emulation of neuronal dynamics such as metastable leakage terms and spiking behaviour (short-term memory), in addition to more stable and persistent changes to the device state (long-term memory). The volatility has an impact on both the equilibrium point and the rate of attainment of the equilibrium point (the time constant of decay). The impact of the volatility state variable on the equilibrium point for a titanium dioxide model was shown in Section III-F1. For stable memristors, the interplay between changes to both rates of switching in the same direction, caused by volatility, and (biased) changes in opposite direction caused by the application of voltage pulses of a given magnitude and time, can lead to memory effects, where states exhibit a dependence on the relative timing of the pulses. A sequence of identical voltage pulses can result in a very different final metastable resistance depending on their spacing in time. For example, in the context of Joule heating, if the voltage pulses are short enough for adiabatic cooling (or relaxation) to occur, the volatility may not increase sufficiently to result in significant switching behaviour. If, however, voltage pulses are more closely spaced, this can result in a larger attained volatility during their application, resulting in switching. Thus, changes in the memristive state are not only dependent on the absolute energy supplied, but also on the way in which this energy is supplied through the pulse shapes and durations.
IV-C Neuromorphic Simulations
@@
We here perform simulations demonstrating the impact of the volatility state variable on two classes of memristor. We first demonstrate the possibility of frequency-dependent switching for stable memristors, and then demonstrate the generation of action potentials for unstable memristors.
IV-C1 Frequency-Dependent Potentiation
@@
Frequency-dependent potentiation has been demonstrated in diffusive memristive devices. Diffusive memristors operate on the basis of the dispersal and regrouping of metal ion clusters under voltage bias [13]. Frequency of application of voltage pulses has been shown to have an effect on the number of pulses required to turn a device ON in protein nanowire catalysed diffusive memristors [25]. This effect was attributed to the resistance drift, due to diffusive regrouping of $Ag+$ ions in the Electrochemical Metallization (ECM) cell, back to the high resistance state during periods where there was no applied bias, meaning that higher frequency voltage pulses required fewer attempts on average to switch devices over a threshold. The number of pulses (of equal magnitude and duration) required to switch the devices from high resistive states to low resistive states was significantly reduced when the frequency of application was increased from 50Hz to 900Hz. A close to linear trend in the number of pulses required as a function of frequency was observed for frequencies higher than 200Hz, with increasing frequency reducing the number of total pulses required to initiate switching. However, a non-linear trend was observed for sub-100Hz frequencies, where the number of pulses increased rapidly. This could be modelled by a volatility state variable whose time constant of decay was on the order of $1/100Hz=0.01s$. In the context of our event-based model, a device with a moderately stable switching characteristic (whose switching rate declines rapidly following the removal of applied bias) can exhibit such frequency dependent behaviour. This is dictated by the state the device is in, as well as the values $V_{a}$ and $V_{\text{off}}$. In Figure 12, we show how changes to the frequency of programming pulses can result in different degrees of potentiation in memristors due to the aforementioned interplay between the volatility and switching. For these experiments, we use a simulation duration of $T=100s$. We update the volatility state variable ($\rho(t)$), at intervals of 0.1s, with updates being generated according to the scheme described in Section II-D3. In these experiments, we supply the same amount of energy to the devices (the same magnitude of voltage over the same amount of time in total), but demonstrate that if the impact of the volatility state variable is high enough, subsequent state transitions can differ greatly, dependent on the frequency of programming, due to the non-linear acceleration in switching effected by an increase in the volatility state variable when the pulses are closely spaced, as compared to sparse pulses in time. We apply square waves of period $T_{\text{pulse}}=1/f$, where $f$ is the frequency of the applied signal. Each square wave period contains a pulse of length 0.1s, and we supply 5 periods as an input signal in each experiment. As such, the same voltage magnitude is applied for the same total amount of time in each case. We modify the parameter $c_{\text{volatile}}=500$, keeping all other parameters as in Table II.
@@ IV-C2 Spiking Behaviour
For unstable memristors, we explore a possible phenomenon enabled by the presence of volatility in a memristive system: the emulation of action potentials.
A “spike” can be defined in the context of a neuromorphic computing architecture as a short-lived increase (or decrease) in an electronic parameter that can be used to process information, such as the voltage, the current, or the conductance. A spike can be thought of as a self-limiting process: once a spike is initiated, it will effect its increase or decrease in the parameter, an associated effect of which is to cause the increase or decrease to cease after a short time. We can envision the creation of spiking behaviour in real devices, if we are able to engineer such a self-limiting process.
We here give a demonstration of how the principle of volatility can be used in the context of the proposed model to generate such a spike. Let us assume that the circuit consists of a memristor in series with a resistor, with an input signal applied across both the memristor and the resistor. This forms a voltage divider circuit, where the proportion of the applied voltage dropped across the memristor will be proportional to its effective instantaneous resistance. If the device is at a high resistance, the proportion of voltage dropped across the memristor will be large. Assuming a volatility whose equilibrium value is proportional to the voltage dropped across the memristor (for example, one can think of Joule heating), the volatility state variable may increase as a result of the magnitude of the applied bias. An increase in the volatility state variable results in a drop in the equilibrium value of the memristor and an increase in the rate of attainment of this value (the net rate). As a result, the device resistance may drop to match the equilibrium point. However, as the resistance decreases, the voltage drops across the memristor, resulting in a decrease in the volatility, which in turn restores the memristor’s equilibrium point to its initial high resistance value, causing the device to return to the high resistance state. One might potentially use the current output of the circuit in order to read and pass on this spike to other parts of the neuromorphic circuit.
As we can see from Equation 10, the equilibrium state is a function of the volatility. Regardless of the parameters, an increase in the volatility results in a shift in the equilibrium state towards the point $n_{eq}=N/2$, as $\lambda_{\text{up}}$ becomes equal to $\lambda_{\text{down}}$. Since the magnitude of both the up and down switching rates also goes up, the rate of attainment of the equilibrium point also goes up. This is visualised in Figure 11.
@@
Biological neurons exhibit two responses to constant input stimuli: firing latencies or rates that are proportional to the magnitude of the input signal (time and frequency encoding, respectively), and a firing response that diminishes over time if the input signal remains constant (adaptation). We propose a method for producing such behaviour using the event-based model, through the dynamics of the volatility state variable.
We can imagine a memristive device with a natural equilibrium state which corresponds to a low resistance. At the same time, we can imagine a volatility state variable which increases in response to the power dissipated by the device. At low resistances, the power dissipated in the memristor will be high, while at high resistances, the reverse will be true, if we assume the voltage drop across the device to be constant.
@@
If the device is at a low resistance, the dissipation of power will result in an increase in the volatility, which results in an increase in the equilibrium resistance (a drop in $n_{eq}$). As the device resistance rises in response, the power dissipated through the device drops. Due to the time constant of decay for volatility, the drop in the equilibrium state will not be instantaneous, and the device resistance will continue to rise. Finally, once the volatility decays sufficiently, the equilibrium resistance will restore to close to its original value, and the device will decrease in resistance, completing a “spike”. The number of such events that happen per unit time will be related to the magnitude of the applied signal. This is shown in Figure 13. In order to produce such a spike train, several parameters were modified, as given in Table IV. The behaviour is essentially brought about through a controlled oscillatory (self-inhibiting) - or transient - response.
Parameter Value
$g_{\text{step}}$ $10^{-8}$
$g_{\text{parallel}}$ $10^{-6}$
$N$ 1000000
$n_{\text{thresh}}$ 800000
$V_{a}$ 1.0
$V_{\text{off}}$0.8
TABLE IV: Parameters to produce spiking behaviour as shown in Figure 13. Only those differing from Table II are listed.
@@
Figure 13: Spiking current behaviour produced using model parameters that produce oscillatory behaviour. Spiking behaviour can be produced through appropriate parameter selection. The intensity of the spikes, the number of spikes in the train, and the frequency of the resultant spike train are dependent on the magnitude of the applied input.
@@
This example demonstrates how through consideration of the volatility state variable as a modelling tool, with its impact on the equilibrium state of the memristor and the rate of attainment of that state, we can predict memory-dependent, self-limiting, short-term behaviour such as spiking, on simulated memristors, offering motivation for its implementation using real memristive devices in a neuromorphic circuit, through the tuning of device parameters.
V Conclusion
@@
We have introduced an event-based modelling strategy for memristive devices where time is the main computational resource. In such systems, event-based simulation is essential, and use of a volatility state variable can help to make clear routes to controlling device behaviour through the spacing of input to the memristive device over time, leading to the possibility of the emulation of different forms of neuronal dynamics, dependent on the stability of the memristive device in question. Our framework, based on multiple parallel Poisson processes, addresses errors associated with discrete time approximation in existing models, as well as being natively suited to event-driven neuromorphic applications. We proposed a volatility state variable, to allow for the capture of transient switching dynamics, motivating this from a physical perspective for various memristive device types. We have also presented a resistive drift dataset for resistive drift in titanium dioxide memristors, which we used to obtain parameters for an example of the model applied to a real device, along with proposing a new linear conductance model of drift in the devices, motivated by a model of the connection between the internal state and the observed resistance. We demonstrated the ability of the event-based model to replicate the resistive drift behaviour under zero bias following the development of the Poisson rate parameter model and volatility model for the titanium dioxide memristors, and subsequent parameter fitting. Finally, we examined the modelling framework from the perspective of neuromorphic computing, studying possible requirements for parameters of future devices that may enable frequency-dependent switching, and replication of action potentials using memristive devices. We introduced the notions of stable and unstable memristors, based on their switching rate in low bias conditions, and demonstrated how our proposed modelling approach based on the volatility state variable allows for better understanding and control of the neuromorphic behaviours. Future work can include applying the modelling framework described to additional device types, specifying probabilistic switching models, volatility models, and readout equations.

### v3

@@ VI Neuromorphic Applications
VI-A Stable and Unstable Memristors
@@
We may consider two types of memristor: those which are highly unstable and those that are stable. For the model as presented in Section IV, the stability of the memristor can be characterised by the rate of attainment of its equilibrium point, which is determined in large part by the magnitude of the energy barrier $E_{a}$. Unstable devices may be useful for the emulation of neuronal body dynamics, where the interplay between volatility and the application of a stimulating input (for example, current spikes) can result in self-limiting spiking responses. In this context, resistive drift happening over short timescales is analogous to the leakage of membrane potential. In contrast, stable memristors may be more useful for the emulation of synaptic weights. In this context, we can model a combination of short and long term memory effects: Short Term Potentiation (STP) can arise for the programming of a stable memristor to a particular state while the volatility is high. The volatility may cause the state variable to partially revert to its equilibrium point, but eventually dissipates, resulting in stability of the resistive state ( Long Term Potentiation (LTP)). For stable memristors, resistive drift is analogous to the phenomenon of synaptic weight decay.
VI-B Volatility and its Impact
@@
The inclusion of volatility in a memristive model adds a key characteristic that could prove crucial to the emulation of neuronal dynamics such as metastable leakage terms and spiking behaviour (short-term memory), in addition to more stable and persistent changes to the device state (long-term memory). The volatility has an impact on both the equilibrium point and the rate of attainment of the equilibrium point (the time constant of decay). The impact of the volatility state variable on the equilibrium point for a titanium dioxide model was shown in Section IV-F1. For stable memristors, the interplay between changes to both rates of switching in the same direction, caused by volatility, and (biased) changes in opposite direction caused by the application of voltage pulses of a given magnitude and time, can lead to memory effects, where states exhibit a dependence on the relative timing of the pulses. A sequence of identical voltage pulses can result in a very different final metastable resistance depending on their spacing in time. For example, in the context of Joule heating, if the voltage pulses are short enough for adiabatic cooling (or relaxation) to occur, the volatility may not increase sufficiently to result in significant switching behaviour. If, however, voltage pulses are more closely spaced, this can result in a larger attained volatility during their application, resulting in switching. Thus, changes in the memristive state are not only dependent on the absolute energy supplied, but also on the way in which this energy is supplied through the pulse shapes and durations.
VI-C Neuromorphic Simulations
@@
We here perform simulations demonstrating the impact of the volatility state variable on two classes of memristor. We first demonstrate the possibility of frequency-dependent switching for stable memristors, and then demonstrate the generation of action potentials for unstable memristors. We choose to use the same rate equation (Boltzmann) and readout equation (linear conductance) modelling components proposed for the TiO2 memristor in Section IV, and use the second volatility component proposed in Section IV-D: structural disruption, mediated by either the voltage for frequency-dependent potentiation (Section VI-C1), or the current for spiking behaviour (Section VI-C2).
VI-C1 Frequency-Dependent Potentiation
@@
Frequency-dependent potentiation has been demonstrated in diffusive memristive devices. Diffusive memristors operate on the basis of the dispersal and regrouping of metal ion clusters under voltage bias [14]. Frequency of application of voltage pulses has been shown to have an effect on the number of pulses required to turn a device ON in protein nanowire catalysed diffusive memristors [26]. The number of pulses (of equal magnitude and duration) required to switch the devices from high to low resistive states was significantly reduced when the frequency of application was increased from 50Hz to 900Hz. A close to linear trend in the number of pulses required as a function of frequency was observed for frequencies higher than 200Hz, with increasing frequency reducing the number of total pulses required to initiate switching. However, a non-linear trend was observed for sub-100Hz frequencies, where the number of pulses increased rapidly. This could be modelled by a volatility state variable whose time constant of decay was on the order of $1/100Hz=0.01s$. In the context of our event-based model, a device with a moderately stable switching characteristic (whose switching rate declines rapidly following the removal of applied bias) can exhibit such frequency dependent behaviour. This is dictated by the state the device is in, as well as the values $V_{a}$ and $V_{\text{off}}$. In Figure 14, we show how changes to the frequency of programming pulses can result in different degrees of potentiation in memristors due to the aforementioned interplay between the volatility and switching. For these experiments, we use a simulation duration of $T=100s$. We incorporate discrete timestep events to the volatility state variable ($\rho(t)$), at intervals of $t_{\text{sample}}=0.1s$, with Euler integration used to perform the updates at events. In these experiments, we supply the same amount of energy to the devices (the same magnitude of voltage over the same amount of time in total), but demonstrate that if the impact of the volatility state variable is high enough, subsequent state transitions can differ greatly, dependent on the frequency of programming, due to the non-linear acceleration in switching effected by an increase in the volatility state variable when the pulses are closely spaced, as compared to sparse pulses in time. We apply square waves of period $T_{\text{pulse}}=1/f$, where $f$ is the frequency of the applied signal. Each square wave period contains a pulse of length 0.1s, and we supply 5 periods as an input signal in each experiment. As such, the same voltage magnitude is applied for the same total amount of time in each case. We modify the parameter $c_{\text{volatile}}=500$, keeping all other parameters as in Table II.
@@ VI-C2 Spiking Behaviour
For unstable memristors, we explore a possible phenomenon enabled by the presence of volatility in a memristive system: the emulation of action potentials. Early on, [3] looked at the Hodgkin-Huxley neuron model [21] as an example of a memristive system. There has since been interest in action potential generation using memristors [27].
@@
A “spike” can be defined in the context of a neuromorphic computing architecture as a short-lived increase (or decrease) in an electronic parameter that can be used to process information, such as the voltage, the current, or the conductance. A spike in biological systems is a self-limiting process. We can envision the creation of spiking behaviour in real devices, if we are able to engineer such a process. Biological neurons exhibit two responses to constant input stimuli: firing latencies or rates that are proportional to the magnitude of the input signal (time and frequency encoding, respectively), and a firing response that diminishes over time if the input signal remains constant (adaptation). We propose a method for producing such behaviour using the event-based model, through the dynamics of the volatility state variable.
As we can see from Equation 10, the equilibrium state is a function of the volatility. Regardless of the parameters, an increase in the volatility results in a shift in the equilibrium state towards the point $n_{eq}=N/2$, as $\lambda_{\text{up}}$ becomes equal to $\lambda_{\text{down}}$. Since the magnitude of both the up and down switching rates goes up, the rate of attainment of the equilibrium point also goes up. This is visualised in Figure 13.
@@
We can imagine a memristive device with a low resistance equilibrium state at rest. Simultaneously, we can imagine a volatility state variable which increases according to the power dissipated by the device. At low resistances, the power dissipated in the memristor will be high, while at high resistances, the reverse will be true. We envision the memristor to be in series with a resistor, forming a voltage divider circuit. At a low resistance, the dissipation of power will result in an increase in the volatility, resulting in an increase in the equilibrium resistance (a drop in $n_{eq}$). As the device resistance rises in response, the power dissipated drops. Due to the volatility time constant, the drop in the equilibrium state will not be instantaneous, and the device resistance will continue to rise. Finally, once the volatility decays sufficiently, the equilibrium resistance will approach its original value, and the resistance will decrease, completing a “spike”. The number of such events that happen per unit time will be related to the magnitude of the applied signal. We model this scenario, using a current-dependent volatility, in Figure 15. To produce such a spike train, several parameters were modified, as given in Table III. The behaviour is essentially brought about through a controlled oscillatory (self-inhibiting) - or transient - response. We note the behaviour resembles spiking dynamics in other complex systems that may emerge in the presence of feedback, such as relaxation oscillations in solid state lasers [28].
@@
Figure 15: Spiking current behaviour produced using model parameters that produce oscillatory behaviour. The controlled oscillations can be mapped to discrete spikes through a thresholding operation. The intensity of the spikes, the number of spikes in the train, and the frequency of the resultant spike train are dependent on the magnitude of the applied input.
Parameter $g_{\text{step}}$ $g_{\text{parallel}}$ NN $n_{\text{thresh}}$ $V_{a}$ $V_{\text{off}}$
Value $10^{-8}$ $10^{-6}$ 1000000 800000 1.0 -0.8
TABLE III: Parameters to produce spiking behaviour as shown in Figure 15. Only those differing from Table II are listed.
@@
This example demonstrates how through consideration of the volatility state variable as a modelling tool, with its impact on the equilibrium state of the memristor and the rate of attainment of that state, we can predict memory-dependent, self-limiting, short-term behaviour such as spiking, on simulated memristors, offering motivation for its implementation using real memristive devices in a neuromorphic circuit, through the tuning of device parameters. Our perspective offers a principled approach for analysing and engineering this behaviour in memristors. Though our proposed neuromorphic examples are hypothetical, they offer guidance for the design of future devices through appropriate parameter selection to tune the relationship between the equilibrium point and its rate of attainment, helping to decode their behaviour and inform their maturation as devices for neuromorphic computing.
VII Conclusion
@@
We have introduced an event-based modelling strategy for memristive devices where time is the main computational resource. In such systems, event-based simulation is essential, and use of a volatility state variable can help to make clear routes to controlling device behaviour through the spacing of input to the memristive device over time, leading to the possibility of the emulation of different forms of neuronal dynamics, dependent on the stability of the memristive device in question. Our framework, based on multiple parallel Poisson processes, addresses errors associated with discrete time approximation in existing models, as well as being natively suited to event-driven neuromorphic applications. We proposed a volatility state variable, to allow for the capture of transient switching dynamics, motivating this from a physical perspective for various memristive device types. We have also presented a resistive drift dataset for resistive drift in titanium dioxide memristors, which we used to obtain parameters for an example of the model applied to a real device, along with proposing a new linear conductance model of drift in the devices as an example of a readout equation, motivated by a model of the connection between the internal state and the observed resistance. We demonstrated the ability of the event-based model to replicate the resistive drift behaviour under zero bias following the development of the Poisson rate parameter model and volatility model for the titanium dioxide memristors, and subsequent parameter fitting. Finally, we examined the modelling framework from the perspective of neuromorphic computing, studying possible requirements for parameters and control strategies for future devices that could enable frequency-dependent switching, and replication of action potentials using memristive devices. Future work can include applying the modelling framework described to additional device types, specifying probabilistic switching models, volatility models, and readout equations.